# A Fractal-based Complex Belief Entropy for Uncertainty Measure in Complex Evidence Theory

### v1

@@ III Fractal-based complex belief entropy
In this section, based on the idea of fractal, a generation process of CPBT is proposed, which can well reveal the distribution of CBBAs of multi-element focal elements. Then FCB entropy is proposed based on the generation process of CPBT, and the maximum entropy model of FCB is deduced.
III-A Theoretical analysis of CPBT from fractal perspective
Compared with the traditional probability theory, DSET and CET can express more information because the mass function is assigned to all power sets in FoD. However, in the real world, the probability of a certain event is usually expressed by the probability theory of Bayesian distribution. Therefore, transforming the mass function of evidence theory into probability is a crucial issue. In DSET, PPT is widely used in practical problems as an important method to convert BBA into probability. In [60], a fractal-based PPT generation process is proposed, which provides a new perspective for describing the relationship between BBA and probability. In CET, the research on the relationship between CBBA and probability has not been widely carried out. CPBT is a common method to convert CBBAs of multi-element focal elements into singletons. Then $\mathbb{C}om$ is used to play the role of decision instead of probability in CET [61]. Inspired by [60], this paper proposes a fractal generation process of CPBT in order to fill the gap in the research of this direction in CET.
For the 2-element FoD $X=\{{{x}_{1}},{{x}_{2}}\}$, the generation process of CPBT can be understood as that the CBBA containing two elements is gradually allocated to a single element in $n$ iterations. When $n$ tends to infinity, the CBBA containing two elements gradually tends to zero, while the CBBA of singletons gradually tends to be stable, and finally the information contained in it tends to be constant. A specific example is given below to understand the generation process of CPBT.
Example III.1.
@@ III-A Theoretical analysis of CPBT from fractal perspective
It is not difficult to understand the proposed CPBT generation process from Example III.1. For FoD with $n$ elements, give a CBBA $\mathbb{M}$. The production process of CPBT based on fractal is that the CBBA value of multi-element focal elements are evenly distributed to single element focal elements in a certain proportion in the process of time change, gradually reducing the influence of non-specificity until finally disappearing. Finally, the CBBA of Bayesian distribution is obtained.
An important feature of fractal is self-similarity. The proposed CPBT generation process is a self-similar iterative process. The fractal diagram of Example 1 is shown in Figure 2.
III-B Fractal-based complex belief entropy
Based on the above CPBT generation process, a new belief entropy called fractal-based complex belief (FCB) entropy is proposed to measure the uncertainty of CBBA. In the above two examples, when the parameters take different values, the size of multi-element focal element allocation is different. In FCB entropy, in order to better reflect the rationality of allocation, it is specified that the focal element is evenly allocated to each of its elements. The calculation diagram of FCB entropy is shown in Figure 3.
@@ III-B Fractal-based complex belief entropy
Figure 3: Flow chart of FCB entropy’s measurement.
Definition III.1.
@@
$\small{{\mathbb{M}}_{F}}({{A}_{k}})=\frac{\mathbb{M}({{A}_{k}})}{{{2}^{\left|{{A}_{k}}\right|}}-1}+\sum\limits_{{{A}_{k}}\subseteq{{B}_{k}}\cap\left|{{A}_{k}}\right|<\left|{{B}_{k}}\right|}{\frac{\mathbb{M}({{B}_{k}})}{{{2}^{\left|{{B}_{k}}\right|}}-1}}.$ (21)
@@
$\small\mathbb{C}o{{m}_{F}}({{A}_{k}})=\frac{\left|{{\mathbb{M}}_{F}}({{A}_{k}})\right|}{\sum\limits_{B\in\Theta}{\left|{{\mathbb{M}}_{F}}({{B}_{k}})\right|}},$ (23)
@@
$\small\sum\limits_{{{A}_{k}}\in{{2}^{\Theta}}}{\mathbb{C}o{{m}_{F}}({{A}_{k}})}=1.$ (24)
@@
$\small{{\mathbb{E}}_{FCB}}(\mathbb{M})=-\sum\limits_{{{A}_{k}}\in{{2}^{\Theta}}}{\frac{\left|{{\mathbb{M}}_{F}}({{A}_{k}})\right|}{\sum\limits_{B\in\Theta}{\left|{{\mathbb{M}}_{F}}({{B}_{k}})\right|}}\log\left(\frac{\left|{{\mathbb{M}}_{F}}({{A}_{k}})\right|}{\sum\limits_{B\in\Theta}{\left|{{\mathbb{M}}_{F}}({{B}_{k}})\right|}}\right)}.$ (25)
@@
$\small\theta\left({{X}_{i}}\right)=\left\{\begin{matrix}\arctan\left(\frac{v}{u}\right),&u>0\\ \frac{\pi}{2},&u=0,v>0\\ -\frac{\pi}{2},&u=0,v<0\\ \arctan\left(\frac{v}{u}\right)+\pi,&u<0,v\geq 0\\ \arctan\left(\frac{v}{u}\right)-\pi,&u<0,v<0\\ \end{matrix}\right.$ (29)
@@
$\small\mathbb{M}({{X}_{i}})=\mathbf{m}({{X}_{i}}){{e}^{i\theta({{X}_{i}})}},$ (30)
$\small\mathbb{M}({{X}_{j}})=\mathbf{m}({{X}_{j}}){{e}^{i\theta({{X}_{j}})}},$ (31)
@@
$\small\mathbb{M}({{X}_{i}})+\mathbb{M}({{X}_{j}})=\mathbf{m}({{X}_{i}}){{e}^{i\theta({{X}_{i}})}}+\mathbf{m}({{X}_{j}}){{e}^{i\theta({{X}_{j}})}}.$ (32)
@@
Through analysis, it can be seen that when $\left|X({{\theta}_{i}})-X({{\theta}_{j}})\right|\in[0,\frac{\pi}{2})\cup(\frac{3\pi}{2},2\pi]$, that is $\cos(X({{\theta}_{i}})-X({{\theta}_{j}}))>0$, $IE(X)>0$, which results in positive interference effect; when $\left|X({{\theta}_{i}})-X({{\theta}_{j}})\right|\in(\frac{\pi}{2},\frac{3\pi}{2})$, that is $\cos(X({{\theta}_{i}})-X({{\theta}_{j}}))<0$, $IE(X)<0$, which results in negative interference effect; when $\left|X({{\theta}_{i}})-X({{\theta}_{j}})\right|=\frac{\pi}{2}$ or $\left|X({{\theta}_{i}})-X({{\theta}_{j}})\right|=\frac{3\pi}{2}$, that is $\cos(X({{\theta}_{i}})-X({{\theta}_{j}}))=0$, $IE(X)=0$, which results in no interference effect. The results of the analysis are shown in the three subgraphs in Figure 4 below.
(a) Positive interference
(b) Null interference
(c) Negative interference
Figure 4: Interference effect between two elements.
The existence of $IE({B}_{k})$ will have different impacts on $\mathbb{C}o{{m}_{F}}({{A}_{k}})$, mainly in the following three cases.
Case 1. For $\left|{{B}_{k}}\right|>\left|{{A}_{k}}\right|$, when $\mathbb{M}({{B}_{k}})$ is evenly distributed to the set $\mathbb{M}({{A}_{k}})$, its complex value and the required complex value tend to the same direction, resulting in a greater degree of support for ${{A}_{k}}$ than the normal state, making $\mathbb{C}o{{m}_{F}}({{A}_{k}})$ larger.
Case 2. For $\left|{{B}_{k}}\right|<\left|{{A}_{k}}\right|$, when $\mathbb{M}({{B}_{k}})$ is evenly distributed to the set $M({{A}_{k}})$, its complex value deviates from the direction of the required complex value of $\mathbb{M}({{A}_{k}})$, resulting in less support than the normal state, making $\mathbb{C}o{{m}_{F}}({{A}_{k}})$ smaller.
Case 3. For $\left|{{B}_{k}}\right|=\left|{{A}_{k}}\right|$, when $\mathbb{M}({{B}_{k}})$ is evenly distributed to the set $\mathbb{M}({{A}_{k}})$, its complex value is consistent with the direction of the required complex value of $\mathbb{M}({{A}_{k}})$, resulting in a support level equal to the normal state, which has no impact on $\mathbb{C}o{{m}_{F}}({{A}_{k}})$.
Axiom III.1.
When CBBA degenerates to BBA, FCB entropy degenerates to FB entropy, that is, In DSET, FCB entropy and FB entropy are equivalent.
Proof III.1.
When CBBA degenerates to BBA, the following formula can be obtained:
$\small{{\mathbf{m}}_{F}}({{A}_{k}})=\frac{\mathbf{m}({{A}_{k}})}{{{2}^{\left|{{A}_{k}}\right|}}-1}+\sum\limits_{{{A}_{k}}\subseteq{{B}_{k}}\cap\left|{{A}_{k}}\right|<\left|{{B}_{k}}\right|}{\frac{\mathbf{m}({{B}_{k}})}{{{2}^{\left|{{B}_{k}}\right|}}-1}}.$ (34)
Because BBA is defined in the real number field, there is no interference effect, namely, $\sum\limits_{{{A}_{k}}\subseteq{{B}_{k}}\cap\left|{{A}_{k}}\right|<\left|{{B}_{k}}\right|}{IE(B_{k})=0}$.
$\small\left|{{\mathbf{m}}_{F}}({{A}_{k}})\right|={{\mathbf{m}}_{F}}({{A}_{k}}),$ (35)
since the value range of BBA is [0,1],
$\small\mathbb{C}o{{m}_{F}}({{A}_{k}})=\frac{\left|{{\mathbf{m}}_{F}}({{A}_{k}})\right|}{\sum\limits_{B\in\Theta}{\left|{{\mathbf{m}}_{F}}({{B}_{k}})\right|}}={{\mathbf{m}}_{F}}({{A}_{k}}).$ (36)
Then simplify the FCB entropy to obtain
${{\mathbb{E}}_{FCB}}(m)={{E}_{FB}}(m)$ $=-\sum\limits_{{{A}_{k}}\in{{2}^{\Theta}}}{\mathbb{C}o{{m}_{F}}({{A}_{k}})\log(\mathbb{C}o{{m}_{F}}({{A}_{k}}))}$ (37)
$=-\sum\limits_{{{A}_{k}}\in{{2}^{\Theta}}}{{{m}_{F}}({{A}_{k}})\log({{m}_{F}}({{A}_{k}}))}.$
Therefore, FCB entropy can be regarded as the generalization of FB entropy, which has better ability to process complex information than FB entropy.
III-C Maximum fractal-based complex belief entropy
@@
$\small\mathbb{E}_{FCB}^{\max}(\mathbb{M})=\log({{2}^{\left|\Theta\right|}}-1)=\log({{2}^{n}}-1),$ (38)
when $\mathbb{C}o{{m}_{F}}({{A}_{k}})=\frac{1}{{{2}^{\left|\Theta\right|}}-1}$.
Proof III.2.
First, FCB entropy is expressed as a function
$\small{{f}_{FCB}}({{A}_{1}},\cdots,{{A}_{{{2}^{\left|\Theta\right|}}-1}})=-\sum\limits_{{{A}_{k}}\in{{2}^{\Theta}}}{\mathbb{C}o{{m}_{F}}({{A}_{k}})\log(\mathbb{C}o{{m}_{F}}({{A}_{k}}))},$ (39)
according to (24), that is, $\sum\limits_{{{A}_{k}}\in{{2}^{\Theta}}}{\mathbb{C}o{{m}_{F}}({{A}_{k}})}=1$. Therefore, the Lagrange function can be constructed as
${{\mathbb{F}}_{FCB}}({{A}_{1}},{{A}_{2}},\cdots,{{A}_{{{2}^{\left|\Theta\right|}}-1}},\lambda)$ $={{f}_{FCB}}({{A}_{1}},{{A}_{2}},\cdots,{{A}_{{{2}^{\left|\Theta\right|}}-1}})$ (40)
$+\lambda(\sum\limits_{{{A}_{k}}\in{{2}^{\Theta}}}{\mathbb{C}o{{m}_{F}}({{A}_{k}})}-1),$
then calculate the first partial derivative of ${{\mathbb{F}}_{FCB}}({{A}_{1}},{{A}_{2}},\cdots,{{A}_{{{2}^{\left|\Theta\right|}}-1}})$. For all $\mathbb{C}o{{m}_{F}}({{A}_{k}})$,
$\small\log(\mathbb{C}o{{m}_{F}}({{A}_{k}}))=\lambda-\frac{1}{\ln a}=k,$ (41)
then according to (24),
$\small\sum\limits_{{{A}_{k}}\in{{2}^{\Theta}}}{\mathbb{C}o{{m}_{F}}({{A}_{k}})}=({{2}^{\left|\Theta\right|}}-1)\mathbb{C}o{{m}_{F}}({{A}_{k}})=1,$ (42)
so when $\mathbb{C}o{{m}_{F}}({{A}_{k}})=\frac{1}{({{2}^{\left|\Theta\right|}}-1)}$, FCB entropy reaches its maximum,
$\small\mathbb{E}_{CFB}^{\max}(\mathbb{M})=-\sum\limits_{{{A}_{k}}\in{{2}^{\Theta}}}{\frac{1}{({{2}^{\left|\Theta\right|}}-1)}\log\frac{1}{({{2}^{\left|\Theta\right|}}-1)}}=\log({{2}^{\left|\Theta\right|}}-1).$ (43)
FB entropy is a generalization of Shannon entropy. Its maximum entropy model is similar to Shannon entropy, and can also deal with decision-making problems in the objective world. As the generalization of FB entropy, the maximum entropy model of FCB entropy has the physical meaning of FB entropy and has more advantages.
@@ IV The properties of the proposed FCB entropy
For the total uncertainty measurement of BBA, different methods generally have 10 properties, which are often used for the analysis of uncertain methods. For the analysis of FCB entropy, the 10 properties have been extended to CET to measure the feasibility and applicability of FCB entropy. The properties of some uncertain methods are summarized in Table I.
@@
Property IV.1 (Probabilistic consistency).
@@
Case 1. when CBBA degenerates BBA and according to Axiom III.1, FCB entropy is degenerated to FB entropy. And the BBA in FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$ satisfies $\sum\limits_{i=1}^{n}{m({{e}_{i}})}=1$, substitute it into (23):
@@
It is obvious that it satisfies Property IV.1.
@@
Property IV.2 (Set correlation).
The number of elements in the focal element will also affect the information modeling when the complex mass function is allocated. The traditional set consistency is controversial, but it provides a new perspective for the study of the set-related properties of FCB entropy. Explanation: Hartley’s information formula measures the amount of information in probability theory, while FCB entropy measures the uncertainty in FoD. In the case of the same dimension, evidence theory can express more information than probability theory. For a FoD with $n$ elements, the framework of evidence theory contains ${{2}^{n}}$ focal elements, while probability theory only contains $n$ elements. Therefore, it is reasonable that the maximum value of FCB entropy is greater than the maximum value of Shannon entropy. For some commonly used entropies, their maximum entropies are summarized in Table II. The maximum value of different entropy changes with the number of elements as shown in Figure 5.

### v2

@@ III Fractal-based complex belief entropy
In this section, the importance of CET in uncertainty modeling is discussed in detail at first. Then we introduce a fractal-inspired generation process for Complex Pignistic Belief Transformation (CPBT) that effectively elucidates the distribution of CBBAs with multi-element focal elements. Subsequently, we propose Fractal-Based (FCB) entropy, which is grounded in the CPBT generation process, and deduce the maximum entropy model for FCB.
III-A The importance of Deng entropy and CET in modeling uncertainty
The classical Dempster-Shafer (D-S) evidence theory, grounded in the real number field, adeptly addresses general decision-making problems [71]. Deng entropy, as an important uncertainty measurement method in D-S evidence theory, has attracted more and more attention [72, 60] and has been improved from multiple perspectives [70, 73, 74] and applied in different fields [75]. Deng entropy takes the effect of cardinality of focal elements in BBA into account in the measurement of uncertainty, which is an important theoretical basis for our FCB entropy [76].
However, the intricacies of real-world scenarios necessitate a more comprehensive representation of information. The use of complex numbers captures the nuances that real numbers alone might not fully express. Traditional D-S evidence theory’s representation is confined to real space, which may fall short in effectively capturing and processing uncertain information, especially when dealing with data that exhibit phase angle information. CBBA offers a more sophisticated and holistic approach to encapsulate all pertinent information about a BBA. Furthermore, while it is true that CBBAs need to be transformed into probabilities for decision-making purposes, this transformation is not straightforward due to the complex nature of the information. The complex mass function must be converted into a form that can be compared and aggregated, which is inherently possible with probabilities. The transformation process itself benefits from the richness of the CBBA, as it provides a more detailed and comprehensive basis for the decision-making process.
Scholarly work on CET’s application in uncertainty modeling is extensive. A negation method for CBBA in CET has been introduced and applied in multi-source information fusion [40, 39]. Additionally, a complex Gaussian fuzzy numbers-based uncertainty modeling technique within CET has facilitated the generation of CBBA, enhancing the precision of post-fusion decision-making beyond traditional approaches [38]. A negation-based multisource information fusion outperforms well-known related approaches with higher classification accuracy and robustness [40].
III-B Theoretical Analysis of CPBT from a Fractal Perspective
DSET and CET offer richer information representation compared to traditional probability theory, as they assign mass functions to the entire power set in the Frame of Discernment (FoD). However, in practical scenarios, the probability of specific events is often articulated through Bayesian probability theory. Thus, the transformation of evidence theory’s mass function into probability is of pivotal importance. In DSET, PPT is frequently employed to convert BBA into probability. [61] introduces a fractal-based PPT generation process, offering a novel lens to understand the connection between BBA and probability. In CET, the translation of CBBA into probability remains relatively unexplored. CPBT is a standard approach for transforming multi-element focal CBBAs into singularities. The decision-making role in CET is then assumed by $\mathbb{C}om$, which replaces the function of probability [40]. Motivated by [61], this paper presents a fractal-based CPBT generation process to address the research gap in this area within CET.
For the 2-element Frame of Discernment (FoD) $X=\{{{x}_{1}},{{x}_{2}}\}$, the Complex Pignistic Belief Transformation (CPBT) generation process can be conceptualized as the gradual allocation of a two-element CBBA to a single element over the course of $n$ iterations. As $n$ approaches infinity, the CBBA associated with the two elements progressively diminishes to zero, while the CBBA of singletons stabilizes, with the information content ultimately converging to a constant value. A detailed example is provided below to elucidate the CPBT generation process.
Example 1.
@@ III-B Theoretical Analysis of CPBT from a Fractal Perspective
It is not difficult to understand the proposed CPBT generation process from Example 1. For FoD with $n$ elements, give a CBBA $\mathbb{M}$. The production process of CPBT based on fractal is that the CBBA value of multi-element focal elements are evenly distributed to single element focal elements in a certain proportion in the process of time change, gradually reducing the influence of non-specificity until finally disappearing. Finally, the CBBA of Bayesian distribution is obtained.
@@
A hallmark characteristic of fractals is their self-similarity, and the proposed CPBT generation process embodies this through a self-similar iterative approach. The fractal diagram illustrating Example 1 is depicted in Figure 2.
III-C Fractal-based complex belief entropy
Drawing on the aforementioned CPBT generation process, we introduce a novel measure of uncertainty for CBBA known as Fractal-Based Complex Belief (FCB) entropy. As demonstrated in the two examples, varying parameter values result in different extents of allocation to multi-element focal elements. To enhance the representation of allocation rationality within FCB entropy, it is stipulated that the focal element’s mass is uniformly distributed among its constituent elements.
Definition 19.
@@ III-C Fractal-based complex belief entropy
${{\mathbb{M}}_{F}}({{A}_{k}})=\frac{\mathbb{M}({{A}_{k}})}{{{2}^{\left|{{A}_{k}}\right|}}-1}+\sum\limits_{{{A}_{k}}\subseteq{{B}_{k}}\cap\left|{{A}_{k}}\right|<\left|{{B}_{k}}\right|}{\frac{\mathbb{M}({{B}_{k}})}{{{2}^{\left|{{B}_{k}}\right|}}-1}}.$ (27)
@@
$\mathbb{C}o{{m}_{F}}({{A}_{k}})=\frac{\left|{{\mathbb{M}}_{F}}({{A}_{k}})\right|}{\sum\limits_{B\in\Theta}{\left|{{\mathbb{M}}_{F}}({{B}_{k}})\right|}},$ (29)
@@
$\sum\limits_{{{A}_{k}}\in{{2}^{\Theta}}}{\mathbb{C}o{{m}_{F}}({{A}_{k}})}=1.$ (30)
@@
${{\mathbb{E}}_{FCB}}(\mathbb{M})=-\sum\limits_{{{A}_{k}}\in{{2}^{\Theta}}}{\frac{\left|{{\mathbb{M}}_{F}}({{A}_{k}})\right|}{\sum\limits_{B\in\Theta}{\left|{{\mathbb{M}}_{F}}({{B}_{k}})\right|}}\log\left(\frac{\left|{{\mathbb{M}}_{F}}({{A}_{k}})\right|}{\sum\limits_{B\in\Theta}{\left|{{\mathbb{M}}_{F}}({{B}_{k}})\right|}}\right)}.$ (31)
@@
$\theta\left({{X}_{i}}\right)=\left\{\begin{matrix}\arctan\left(\frac{v}{u}\right),&u>0\\ \frac{\pi}{2},&u=0,v>0\\ -\frac{\pi}{2},&u=0,v<0\\ \arctan\left(\frac{v}{u}\right)+\pi,&u<0,v\geq 0\\ \arctan\left(\frac{v}{u}\right)-\pi,&u<0,v<0\\ \end{matrix}\right.$ (35)
@@
$\mathbb{M}({{X}_{i}})=\mathbf{m}({{X}_{i}}){{e}^{i\theta({{X}_{i}})}},$ (36)
$\mathbb{M}({{X}_{j}})=\mathbf{m}({{X}_{j}}){{e}^{i\theta({{X}_{j}})}},$ (37)
@@
$\mathbb{M}({{X}_{i}})+\mathbb{M}({{X}_{j}})=\mathbf{m}({{X}_{i}}){{e}^{i\theta({{X}_{i}})}}+\mathbf{m}({{X}_{j}}){{e}^{i\theta({{X}_{j}})}}.$ (38)
@@
Axiom 1.
When CBBA degenerates to BBA, FCB entropy degenerates to FB entropy, that is, In DSET, FCB entropy and FB entropy are equivalent. The proof 1 is in the appendix.
III-D Maximum fractal-based complex belief entropy
@@
$\mathbb{E}_{FCB}^{\max}(\mathbb{M})=\log({{2}^{\left|\Theta\right|}}-1)=\log({{2}^{n}}-1),$ (40)
when $\mathbb{C}o{{m}_{F}}({{A}_{k}})=\frac{1}{{{2}^{\left|\Theta\right|}}-1}$. The proof 2 is in the appendix.
FCB entropy represents an extension of Shannon entropy, with its maximum entropy model paralleling that of Shannon’s while also being capable of addressing decision-making issues in the real world. As an advancement over FB entropy, the maximum entropy model of FCB entropy retains the physical interpretation of FB entropy and offers additional advantages.
@@ IV The properties of the proposed FCB entropy
For the total uncertainty measurement of BBA, different methods generally have 10 properties [77], which are often used for the analysis of uncertain methods. For the analysis of FCB entropy, the 10 properties have been extended to CET to measure the feasibility and applicability of FCB entropy. The properties of some uncertain methods are summarized in Table I.
@@
Property 1 (Probabilistic consistency).
@@
Case 1. when CBBA degenerates BBA and according to Axiom 1, FCB entropy is degenerated to FB entropy. And the BBA in FoD $\Theta=\{{{e}_{1}},{{e}_{2}},\cdots,{{e}_{n}}\}$ satisfies $\sum\limits_{i=1}^{n}{m({{e}_{i}})}=1$, substitute it into (29):
@@
It is obvious that it satisfies Property 1.
@@
Property 2 (Set correlation).
The number of elements in the focal element will also affect the information modeling when the complex mass function is allocated. The traditional set consistency is controversial, but it provides a new perspective for the study of the set-related properties of FCB entropy.
Explanation: Hartley’s information formula measures the amount of information in probability theory, while FCB entropy measures the uncertainty in FoD. In the case of the same dimension, evidence theory can express more information than probability theory. For a FoD with $n$ elements, the framework of evidence theory contains ${{2}^{n}}$ focal elements, while probability theory only contains $n$ elements. Therefore, it is reasonable that the maximum value of FCB entropy is greater than the maximum value of Shannon entropy. For some commonly used entropies, their maximum entropies are summarized in Table II. The maximum value of different entropy changes with the number of elements as shown in Figure 3.